# Automatic gradient descent with generalized Newton’s method

### v1

@@ Abstract
We propose the generalized Newton’s method (GeN) — a Hessian-informed approach that applies to any optimizer such as SGD and Adam, and covers the Newton-Raphson method as a sub-case. Our method automatically and dynamically selects the learning rate that accelerates the convergence, without the intensive tuning of the learning rate scheduler. In practice, out method is easily implementable, since it only requires additional forward passes with almost zero computational overhead (in terms of training time and memory cost), if the overhead is amortized over many iterations. We present extensive experiments on language and vision tasks (e.g. GPT and ResNet) to showcase that GeN optimizers match the state-of-the-art performance, which was achieved with carefully tuned learning rate schedulers. Code to be released at https://github.com/ShiyunXu/AutoGeN.
@@ 1 Introduction
BFGS [9, 20, 60], LBFGS [11] need to tune $\approx\mathbf{H}$ $d^{2}$
GeN (ours) $={\mathbf{G}^{\top}\mathbf{g}/\mathbf{g}^{\top}\mathbf{H}\mathbf{g}}$ Any $1$ or $d$
D-adaptation [13], Prodigy [46], DoG [29], DoWG [32] $\approx{\|\bm{w}_{0}-\bm{w}_{*}\|}/{\max\mathbf{G}}$ $=\mathbf{I}$ or $\sqrt{\text{diag}(\mathbf{F})}$ $1$ or $d$
AdaHessian [71], Sophia [37] need to tune $\approx\text{diag}(\mathbf{H})$ $d$
Shampoo [22], K-FAC [44], Natural Gradient [4] need to tune $\approx\mathbf{F}$ $d$
Adam [33], AdamW [41], AdaGrad [18], AdaDelta [73], RMSProp [25] need to tune $\approx\sqrt{\text{diag}(\mathbf{F})}$ $d$
SGD [59], Heavyball [52], NAG [47] need to tune $=\mathbf{I}$ $1$
For large-scale optimization problems with millions to billions of parameters, the training can be significantly costly. For example, large language models are trained with trillions of tokens, on thousands of GPUs, costing millions of dollars [61, 6], and at high carbon footprint [51, 42, 16, 65]: GPT-3 (175B) [8], LLAMA (7$\sim$70B) [65, 66], Chinchilla (70B) [26], PaLM (540B) [12], and Falcon (7$\sim$180B) [3] models are trained on $\gtrapprox 1$T tokens; state-of-the-art models require extremely long training time, measured in thousands of PetaFLOPS-days [3, Figure 2] or millions of GPU hours [2]. It is therefore computationally prohibitive to instantiate a dense $\mathbf{P}_{t}$ or to tune the hyperparameters of $\eta_{t}$. However, the second-order pre-conditioner and the learning rate scheduler are critical to the fast convergence.
@@
On the other hand, manually setting a learning rate scheduler requires a critical balance between the utility (more hyperparameters the better) and the tuning effort (less hyperparameters the easier). The simplest approach is to set a constant learning rate $\eta_{t}=\eta$ (only one hyperparameter) through grid search. However, the mini-batch gradient descent with a constant $\eta_{t}$ may not allow the model to converge even under the convex setting, which is easier to optimize than the non-convex setting in deep learning. Additionally, as we see in Figure 1, small $\eta$ converges slow but to better solution, and vice versa for large $\eta$, which depicts the importance of setting the proper $\eta_{t}$ at each iteration. That is, tuning the learning rate scheduler is essentially selecting $T$ (number of iterations) learning rates to maximize the convergence speed. To accommodate this challenge, a line of schedulers (including the linear decay, warm-up and cosine scheduler) has been proposed to control the learning rate during the training, with a number of hyperparameters. For instance, LLAMA 65B [65] uses a scheduler with 3 hyperparameters: warm-up steps=2000, maximum $\eta_{t}=1.5e^{-4}$ and final $\eta_{t}=1.5e^{-5}$. Unfortunately, most schedulers are heuristic and have their limits: they do not guarantee fast convergence and become increasingly difficult to tune as the model and data sizes increase.
@@
This work is closely related to previous literature in learning rate schedulers (especially the parameter-free methods), optimization with Hessian information, and deep learning system design, which are discussed in Section 3.2 and Appendix D.
@@
We propose GeN – a Hessian-informed approach which merges the information from the full Hessian into the learning rate, so as to dynamically adapt to the loss landscape and accelerate the convergence. We examine GeN on 5 criteria: automaticity, applicability, scalability, computation efficiency, and convergence speed.
To our best knowledge, GeN is the first automatic optimizer that is applicable to the general optimizers [^1] We refer to Related Works in Appendix D for comparison with other heuristic or automatic learning rates.. Our efficiency analysis shows that GeN is perfectly scalable to large-scale training and as computationally efficient as existing optimizers. We empirically demonstrate that GeN is highly performant on image classification, text classification, natural language generation, object detection, instance segmentation, recommendation system, and parameter-efficient training (PET).
@@ 2.2 Second-order Taylor expansion
In deep learning, the models are updated via the gradient descent via the optimizers:
@@
In light of (2.2), we claim that employing the second-order Taylor expansion is (1) necessary, since the first-order Taylor expansion only characterizes a linear approximation of the loss and thus fail to characterize the loss curvature; and (2) sufficient, especially in the small learning rate regime where large models are universally trained [^2] The classical convergence analysis of gradient descent also leverages the second-order Taylor expansion $L(w-\eta\mathbf{G})\leq L(w)+\eta\mathbf{G}^{\top}\mathbf{G}+\frac{\mathcal{L}\eta^{2}}{2}||\mathbf{G}||^{2}$ where $\mathcal{L}$ is the Lipschitz smoothness of $L$. This quadratic function is minimized at $\eta=1/\mathcal{L}$.. In fact, (2.2) is also used in recent works on the neural scaling laws [31, 45].
@@
Figure 2: Illustration of the second-order Taylor expansion in (2.3). Left two: ResNet18 on CIFAR100. Right two: GPT2 on E2E.
@@ 3.1 Generalized Newton’s method
$\eta_{t}^{*}=\mathbf{G}_{t}^{\top}\mathbf{g}_{t}^{\text{optim}}\Big{/}(\mathbf{g}_{t}^{\text{optim}})^{\top}\mathbf{H}_{t}\mathbf{g}_{t}^{\text{optim}}$ (3.1)
@@
$\bm{w}_{t+1}=\bm{w}_{t}-\eta_{t}^{*}\mathbf{g}_{t}^{\text{optim}}=\bm{w}_{t}-\frac{(\mathbf{g}_{t}^{\text{optim}})^{\top}\mathbf{G}_{t}\mathbf{g}_{t}^{\text{optim}}}{(\mathbf{g}_{t}^{\text{optim}})^{\top}\mathbf{H}_{t}\mathbf{g}_{t}^{\text{optim}}}$ (3.2)
@@
$(\mathbf{g}_{t}^{\top}\mathbf{H}_{t})_{R}^{-1}=\mathbf{g}_{t}/\mathbf{g}_{t}^{\top}\mathbf{H}_{t}\mathbf{g}_{t}\Longrightarrow\eta_{t}^{*}\mathbf{g}_{t}=(\mathbf{g}_{t}^{\top}\mathbf{H}_{t})_{R}^{-1}\mathbf{g}_{t}^{\top}\mathbf{G}_{t}$
@@ 3.2 Deriving $\eta^{*}$ without back-propagation
Alternatively, we can estimate $\mathbf{g}_{t}^{\top}\mathbf{H}_{t}\mathbf{g}_{t}$ via the Hessian-vector product of $\mathbf{H}_{t}\mathbf{g}_{t}$, without directly accessing $\mathbf{H}_{t}$: firstly back-propagating on $L$ gives $\mathbf{g}_{t}$ and then back-propagating on $\mathbf{g}_{t}(\bm{w}_{t})^{\top}\mathbf{g}_{t}$ gives $\mathbf{H}_{t}\mathbf{g}_{t}$. In fact, all above-mentioned methods (except BFGS and Fisher-related methods) need the Hessian-vector product [^4] Estimating the diagonal Hessian needs the Hutchinson method and Hessian-vector product, where $\frac{1}{K}\sum_{k\leq K}\mathbf{v}_{k}\odot\mathbf{H}\mathbf{v}_{k}\to\text{diag}(\mathbf{H})$ as $K\to\infty$ for $\mathbf{v}_{k}\sim N(0,I)$. Because the precision of these methods relies on computing $\mathbf{v}_{k}\odot\mathbf{H}\mathbf{v}_{k}$ for $K$ rounds, the training is very slow as the cost is $K$ times that of a standard back-propagation, say $K=20$ in Sophia and $K=100$ in PyHessian., which is not supported in large-scale distributed learning (e.g. DeepSpeed [58], Megatron [62], and FSDP [19]), because the computation graph is complicated and interferes with the communication orchestra. In summary, existing methods suffer from two constraints: they have to either approximate $\mathbf{H}_{t}$ (sacrificing the performance) or rely on the Hessian-vector product (sacrificing the efficiency and scalability).
In stark contrast, we can overcome these constraints by using two additional forward passes to efficiently compute $\mathbf{G}_{t}^{\top}\mathbf{g}_{t}$ and $\mathbf{g}_{t}^{\top}\mathbf{H}_{t}\mathbf{g}_{t}$ without accessing $\mathbf{G}_{t}$ and $\mathbf{H}_{t}$, which is universally easy-to-implement because the forward pass is a fundamental operation in any deep learning system. To be specific, we apply the gradient descent with different learning rates and derive the quadratic function in (2.3) (see Algorithm 1). As an extension, our method can use more forward passes to fit a higher-order polynomial than the second-order one in (2.3).
@@ 3.3 Algorithm
We now present an efficient algorithm that implements (3.2). Specifically, we apply
$\eta^{*}\approx\frac{\eta_{t-1}}{2}\frac{L_{+}-L_{-}}{L_{+}-2L_{0}+L_{-}}$ (3.3)
to approximate the optimal learning rate in (3.1), where $L_{\pm}:=L(\bm{w}_{t}\pm\eta_{t-1}\mathbf{g}_{t}^{\text{optim}})$.
@@
6:         Fit the quadratic function in (2.3):$(-\eta_{t-1},0,\eta_{t-1})\to(L_{-},L_{0},L_{+})$
7:         Derive the optimal learning rate $\eta_{t}=\gamma\eta_{t-1}+(1-\gamma)\eta^{*}$ by (3.3)
8:     Update $\bm{w}_{t+1}=\bm{w}_{t}-\eta_{t}\mathbf{g}_{t}$
@@
  • In line 6, we highlight that $(-\eta_{t-1},0,\eta_{t-1})$ is symmetric and auto-regressive, without introducing a new hyperparameter. The choice is justified in Section A.3. To derive (2.3), we can either fit a parabolic curve or equivalently use the finite difference method.

### v3

# Gradient descent with generalized Newton’s method

@@ Abstract
We propose the generalized Newton’s method (GeN) — a Hessian-informed approach that applies to any optimizer such as SGD and Adam, and covers the Newton-Raphson method as a sub-case. Our method automatically and dynamically selects the learning rate that accelerates the convergence, without the intensive tuning of the learning rate scheduler. In practice, our method is easily implementable, since it only requires additional forward passes with almost zero computational overhead (in terms of training time and memory cost), if the overhead is amortized over many iterations. We present extensive experiments on language and vision tasks (e.g. GPT and ResNet) to showcase that GeN optimizers match the state-of-the-art performance, which was achieved with carefully tuned learning rate schedulers.
@@ 1 Introduction
BFGS (Broyden, 1970), LBFGS (Byrd et al., 1995) need to tune $\approx\mathbf{H}$ $d^{2}$
GeN (this work) $={\mathbf{G}^{\top}\mathbf{g}/\mathbf{g}^{\top}\mathbf{H}\mathbf{g}}$ Any $1$ or $d$
D-adaptation (Defazio & Mishchenko, 2023), Prodigy (Mishchenko & Defazio, ), DoG (Ivgi et al., 2023), DoWG (Khaled et al., 2023) $\approx{\|\bm{w}_{0}-\bm{w}_{*}\|}/{\max\mathbf{G}}$ $=\mathbf{I}$ or $\sqrt{\text{diag}(\mathbf{F})}$ $1$ or $d$
AdaHessian (Yao et al., 2021), Sophia (Liu et al., ) need to tune $\approx\text{diag}(\mathbf{H})$ $d$
Shampoo (Gupta et al., 2018), K-FAC (Martens & Grosse, 2015), Natural Gradient (Amari, 1998) need to tune $\approx\mathbf{F}$ $d$
Adam (Kingma & Ba, 2014), AdamW (Loshchilov & Hutter, 2017), AdaGrad (Duchi et al., 2011), AdaDelta (Zeiler, 2012), RMSProp (Hinton et al., 2012) need to tune $\approx\sqrt{\text{diag}(\mathbf{F})}$ $d$
SGD (Robbins & Monro, 1951), Heavyball (Polyak, 1964), NAG (Nesterov, 1983) need to tune $=\mathbf{I}$ $1$
For large-scale optimization problems with millions to billions of parameters, the training can be significantly costly. For example, large language models are trained with trillions of tokens, on thousands of GPUs, costing millions of dollars (Sharir et al., 2020; Biderman et al., 2023), and at high carbon footprint (Patterson et al., 2021; Luccioni et al., 2023; Dodge et al., 2022; Touvron et al., 2023a): GPT-3 (175B, Brown et al. (2020)), LLAMA (7$\sim$70B, Touvron et al. (2023a; b)), Chinchilla (70B, Hoffmann et al. (2022)), PaLM (540B, Chowdhery et al. (2023)), and Falcon (7$\sim$180B, Almazrouei et al. (2023)) models are trained on $\gtrapprox 1$T tokens; state-of-the-art models require extremely long training time, measured in thousands of PetaFLOPS-days (Almazrouei et al., 2023, Figure 2) or millions of GPU hours (AI@Meta, 2024). It is therefore computationally prohibitive to instantiate a dense $\mathbf{P}_{t}$ or to tune the hyperparameters of $\eta_{t}$. However, the second-order pre-conditioner and the learning rate scheduler are critical to the fast convergence.
@@
On the other hand, manually setting a learning rate scheduler requires a critical balance between the utility — more hyperparameters the better, and the tuning effort — fewer hyperparameters the easier. The simplest approach is to set a constant learning rate $\eta_{t}=\eta$ (only one hyperparameter) through grid search. However, the mini-batch gradient descent with a constant $\eta_{t}$ may not allow the model to converge even under the convex setting, which is easier to optimize than the non-convex setting in deep learning. Additionally, as we see in Figure 1, small $\eta$ converges slow but to better solution, and vice versa for large $\eta$, which depicts the importance of setting the proper $\eta_{t}$ at each iteration. That is, tuning the learning rate scheduler is essentially selecting $T$ (number of iterations) learning rates to maximize the convergence speed. To accommodate this challenge, a line of schedulers (including the linear decay, warm-up and cosine scheduler) has been proposed to control the learning rate during the training, with a number of hyperparameters. For instance, LLAMA 65B (Touvron et al., 2023a) uses a scheduler with 3 hyperparameters: warm-up steps=2000, maximum $\eta_{t}=1.5e^{-4}$ and final $\eta_{t}=1.5e^{-5}$. Unfortunately, most schedulers are heuristic and have their limits: they do not guarantee fast convergence and become increasingly difficult to tune as the model and data sizes increase.
@@
This work is closely related to previous literature in learning rate schedulers (including the heuristic and the parameter-free methods), optimization with Hessian information, and deep learning system design, which are discussed in Section 3.2 and Appendix D.
@@
We propose the generalized Newton’s method (GeN) — a Hessian-informed approach which merges the information from the full Hessian into the learning rate, so as to dynamically adapt to the loss landscape and accelerate the convergence. We examine GeN on 5 criteria: automaticity, applicability, scalability, computation efficiency, and convergence speed.
Overall, GeN [^1] We use ‘generalized’ to mean that (1) Newton-Raphson method can be viewed as a sub-case of GeN and (2) we leverage the generalized inverse without inversing the Hessian matrix. is an automatic optimizer that is applicable to the general optimizers and work successfully in deep learning without adding much computations. Our efficiency analysis shows that GeN is perfectly scalable to large-scale training and as computationally efficient as existing optimizers. We empirically demonstrate that GeN is highly performant on image classification, text classification, natural language generation, object detection, instance segmentation, recommendation system, and parameter-efficient training (PET).
@@ 2.2 Second-order Taylor expansion
In deep learning, the models are updated via the gradient descent by various optimizers:
@@
In light of (2.2), we claim that employing the second-order Taylor expansion is (1) necessary, since the first-order Taylor expansion only characterizes a linear approximation of the loss and thus fail to characterize the loss curvature; and (2) sufficient, especially in the small learning rate regime where large models are universally trained and $O(\eta^{3})$ is negligible [^2] The classical convergence analysis of gradient descent also leverages the second-order Taylor expansion $L(w-\eta\mathbf{G})\leq L(w)+\eta\mathbf{G}^{\top}\mathbf{G}+\frac{\mathcal{L}\eta^{2}}{2}||\mathbf{G}||^{2}$ where $\mathcal{L}$ is the Lipschitz smoothness of $L$. This quadratic function is minimized at $\eta=1/\mathcal{L}$.. In fact, (2.2) is also used in the theoretical analysis of neural scaling laws (Kaplan et al., 2020; McCandlish et al., 2018).
@@
Figure 2: Illustration of the second-order Taylor expansion in (2.3). Left two: ResNet18 on CIFAR100 with SGD. Right two: GPT2 on E2E with AdamW.
@@ 3.1 Generalized Newton’s method
$\eta_{\text{GeN}}^{*}(\mathbf{g}_{t}^{\text{optim}})=\mathbf{G}_{t}^{\top}\mathbf{g}_{t}^{\text{optim}}\Big{/}(\mathbf{g}_{t}^{\text{optim}})^{\top}\mathbf{H}_{t}\mathbf{g}_{t}^{\text{optim}}$ (3.1)
@@
$\bm{w}_{t+1}=\bm{w}_{t}-\eta_{\text{GeN},t}^{*}\mathbf{g}_{t}^{\text{optim}}=\bm{w}_{t}-\frac{(\mathbf{g}_{t}^{\text{optim}})^{\top}\mathbf{G}_{t}\mathbf{g}_{t}^{\text{optim}}}{(\mathbf{g}_{t}^{\text{optim}})^{\top}\mathbf{H}_{t}\mathbf{g}_{t}^{\text{optim}}}$ (3.2)
@@
Remark 3.2.
When the formulation of (3.2) is restricted to SGD without the pre-conditioning, this equivalent form of GeN-SGD has also been presented in prior works like LQA (Zhu et al., 2021) and QLABGrad (Fu & Wu, 2024) and ADLER (Balboni & Bacciu, 2023). However, without the general $\mathbf{g}^{\text{optim}}$ through pre-conditioning, such formulation does not include the Newton’s method as a sub-case. We dedicate Section D.4 to highlight the similarities and differences between GeN and these works.
@@
$(\mathbf{g}_{t}^{\top}\mathbf{H}_{t})_{R}^{-1}=\mathbf{g}_{t}/\mathbf{g}_{t}^{\top}\mathbf{H}_{t}\mathbf{g}_{t}\Longrightarrow\eta_{t}^{*}\mathbf{g}_{t}=(\mathbf{g}_{t}^{\top}\mathbf{H}_{t})_{R}^{-1}\mathbf{g}_{t}^{\top}\mathbf{G}_{t}$ (3.3)
@@ 3.2 Deriving $\eta^{*}$ without back-propagation
Alternatively, we can estimate $\mathbf{g}_{t}^{\top}\mathbf{H}_{t}\mathbf{g}_{t}$ via the Hessian-vector product of $\mathbf{H}_{t}\mathbf{g}_{t}$, without directly accessing $\mathbf{H}_{t}$: firstly back-propagating on $L$ gives $\mathbf{g}_{t}$ and then back-propagating on $\mathbf{g}_{t}(\bm{w}_{t})^{\top}\mathbf{g}_{t}$ gives $\mathbf{H}_{t}\mathbf{g}_{t}$. In fact, all above-mentioned methods (except BFGS and Fisher-related methods) need the Hessian-vector product [^4] Estimating the diagonal Hessian needs the Hutchinson method and Hessian-vector product, where $\frac{1}{K}\sum_{k\leq K}\mathbf{v}_{k}\odot\mathbf{H}\mathbf{v}_{k}\to\text{diag}(\mathbf{H})$ as $K\to\infty$ for $\mathbf{v}_{k}\sim N(0,I)$. Because the precision of these methods relies on computing $\mathbf{v}_{k}\odot\mathbf{H}\mathbf{v}_{k}$ for $K$ rounds, the training is very slow as the cost is $K$ times that of a standard back-propagation, say $K=20$ in Sophia and $K=100$ in PyHessian., which is not supported in large-scale distributed learning such as DeepSpeed (Rasley et al., 2020), Megatron (Shoeybi et al., 2019), and FSDP (FairScale authors, 2021), because the computation graph is complicated and interferes with the communication orchestra. In summary, existing methods suffer from two constraints: they have to either approximate $\mathbf{H}_{t}$ and sacrifice the performance, or to rely on the Hessian-vector product and sacrifice the efficiency and scalability.
In stark contrast, we can overcome these constraints by using multiple additional forward passes to efficiently compute $\mathbf{G}_{t}^{\top}\mathbf{g}_{t}$ and $\mathbf{g}_{t}^{\top}\mathbf{H}_{t}\mathbf{g}_{t}$ without accessing $\mathbf{G}_{t}$ and $\mathbf{H}_{t}$, which is universally easy-to-implement because the forward pass is a fundamental operation in any deep learning system. To be specific, we apply gradient descent with different learning rates and derive the quadratic function in (2.3). Our approach can be viewed as an extension of LQA (Zhu et al., 2021), but allowing more flexibility while enhancing the stability and speed of convergence (see Algorithm 1 and Remark 3.5). As an extension, our method can use more forward passes to fit a higher-order polynomial than the second-order one in (2.3).
@@ 3.3 Algorithm
We now present a quadratic curve fitting approach to estimate the optimal learning rate $\eta_{\text{GeN}}^{*}$ in (3.1):
$\eta^{*}(\Gamma)=b^{*}/A^{*}\text{ where }A^{*},b^{*}=\underset{A\in\mathbb{R},b\in\mathbb{R}}{\text{argmin}}\sum_{\eta\in\Gamma}\left|L(\bm{w}_{t}-\eta\mathbf{g}_{t}^{\text{optim}})-L(\bm{w}_{t})+b\eta-A\frac{\eta^{2}}{2}\right|^{2}.$ (3.4)
in which $\Gamma$ is a set of learning rates, e.g. $\Gamma=\{-\eta_{t-1},0,\eta_{t-1}\}$ in Algorithm 1, and the objective function leverages the Taylor expansion in (2.3). Therefore, we obtain
$\mathbf{G}_{t}^{\top}\mathbf{g}_{t}^{\text{optim}}\approx b^{*}\text{ and }(\mathbf{g}_{t}^{\text{optim}})^{\top}\mathbf{H}_{t}\mathbf{g}_{t}^{\text{optim}}\approx A^{*}\Longrightarrow\eta^{*}\approx\eta^{*}_{\text{GeN}}.$
In fact, a mathematical equivalence can be established between this curve fitting and the finite difference (see Section A.3). Especially, for 3-point $\Gamma=\{-\eta_{t-1},0,\eta_{t-1}\}$, the finite difference approach from Zhu et al. (2021) gives an estimate that is equal to (3.4):
$\eta^{*}_{\text{LQA}}(\Gamma)=\frac{\eta_{t-1}}{2}\frac{L_{+}-L_{-}}{L_{+}-2L_{0}+L_{-}}\text{ where $L_{\pm}:=L(\bm{w}_{t}\pm\eta_{t-1}\mathbf{g}_{t}^{\text{optim}})$}$ (3.5)
However, on one hand, (3.5) is not generalizable as it takes different forms whenever $\Gamma$ changes (see an example of $\eta_{\text{LQA}}^{*}(\{0,\eta,2\eta\})$ in Section A.3). On the other hand, the formulation is complicated when $\Gamma$ takes more than 3 points (see an example of 5 points in (A.2)), while (3.4) easily extends to any number of points as demonstrated in Figure 2.
In practice, we leverage the curve fitting to implement (3.2) via an efficient algorithm in Algorithm 1.
@@
6:         Fit the quadratic function via (3.4): $\{-\eta_{t-1},0,\eta_{t-1}\}\to\{L_{-},L_{0},L_{+}\}$
7:         Extract $A^{*},b^{*}$ from the quadratic function and derive $\eta^{*}=b^{*}/A^{*}$
[…]
9:              Update the learning rate $\eta_{t}=\gamma\eta_{t-1}+(1-\gamma)\eta^{*}$
10:     Update $\bm{w}_{t+1}=\bm{w}_{t}-\eta_{t}\mathbf{g}_{t}$
@@
  • In line 6, we highlight that $\{-\eta_{t-1},0,\eta_{t-1}\}$ is symmetric and auto-regressive, without introducing a new hyperparameter. The choice is justified in Section A.3.
  • In line 8, we ensure the convexity ($A^{*}>0$), the positivity ($b^{*}>0$) and the goodness of fit (R2 score $>0.99$) for the update of learning rate.
  • Algorithm 1 is an approximation to the GeN method in (3.2), since (3.4) approximates $\eta_{\text{GeN}}$ in (3.1). We also give Algorithm 2 that implements GeN exactly as described in Section 3.2.